Drive the expression l=nAevdDrive the expression l=nAevd

Grade:12

1 Answers

Arun
6 years ago
Dear Ankita
 
let no. of electrons per unit volume=n
charge on an atom=e
no.of electrons in length l of the conductor=n×volume of conductor=nAl
total charge contained in the conductor is
q = neAl
all the electrons which enter the conductor at the right end will pass through the left end of the conductor in time,
t = distance/velocity
=l/vd
therefore,current,
I=q/t=enAl/(l/vd)
=> I=neAvd
 
Regards
Arun (askIITians forum expert)